OpenStudy (anonymous):

I need some help with multivariable calculus. Stuck, and would like some clarification. Question in thread.

OpenStudy (anonymous):

It is given that \(g(r,\theta)=f(r\cos \theta, r\sin \theta)\) i need to express \[\frac{\partial ^2 g}{\partial \theta^2}\] in terms of the partial derivatives of f. So let \(x(r,\theta) = r\cos \theta \) and \(y(r,\theta) = r\sin \theta \), thus we have \(f(x(r,\theta),y(r,\theta))\) \[\begin{split} \frac{\partial g}{\partial \theta} &= \frac{\partial f}{\partial x}\frac{\partial x}{\partial \theta} +\frac{\partial f}{\partial y}\frac{\partial y}{\partial \theta} \\ &= \frac{\partial f}{\partial x}\cdot -r\sin\theta+\frac{\partial f}{\partial y}\cdot r\cos\theta\\\\ \end{split} \] Therefore, \[\frac{\partial^2 g }{\partial \theta^2} =\frac{d}{d \theta} \left[ \frac{\partial f}{\partial x}\cdot -r\sin\theta +\frac{\partial f}{\partial y}\cdot r\cos\theta\right] \] This is what i have done so far, but i am not sure how to carry on. I know i would need to use product rule, but i am not sure what \[\frac{d}{d \theta} \left[ \frac{\partial f}{\partial x}\right] \] is. Now Wio did try explaining that \[\begin{split} \frac{d}{d\theta }\left[\frac{\partial f}{\partial x}\frac{\partial x}{\partial \theta}\right] &=\frac{d}{d\theta}\left[\frac{\partial f}{\partial\theta}\right]\frac{\partial x}{\partial \theta}+\frac{\partial f}{\partial x}\frac{d}{d\theta}\left[\frac{\partial x}{\partial \theta}\right] \\ &= \left(\frac{\partial^2 f}{\partial x^2}\frac{\partial x}{\partial \theta}\right)\frac{\partial x}{\partial\theta}+\frac{\partial f}{\partial x}\frac{\partial^2 x}{\partial \theta^2} \\ &=\frac{\partial^2 f}{\partial x^2}\left(\frac{\partial x}{\partial \theta}\right)^2+\frac{\partial f}{\partial x}\frac{\partial^2 x}{\partial \theta^2} \end{split} \] But I do need a bit of clarification about how the term on the RHS \[\frac{d}{d\theta}\left[\frac{\partial f}{\partial\theta}\right]\] came from \[\frac{d}{d\theta }\left[\frac{\partial f}{\partial x}\frac{\partial x}{\partial \theta}\right] \].

OpenStudy (anonymous):

Seems like the df/d(theta) right after the equals sign should be a df/dx

OpenStudy (anonymous):

oh ok, so if thats the case that result still holds, and thats how i should finish off the question.

OpenStudy (anonymous):

Yup looks good

OpenStudy (anonymous):

one quick question why is \[=\frac{d}{d\theta}\left[\frac{\partial f}{\partial x}\right]= \left(\frac{\partial^2 f}{\partial x^2}\frac{\partial x}{\partial \theta}\right)\]

OpenStudy (anonymous):

@saifoo.khan do you think you can help me?

OpenStudy (saifoo.khan):

I'm really sorry. I can't. :S

OpenStudy (anonymous):

have you looked through the whole post? do you know anyone who would be able to explain this to me?

OpenStudy (saifoo.khan):

Many helpers aren't online currently. @jim_thompson5910 might know this.

OpenStudy (anonymous):

ok, well i will leave it open for a while

OpenStudy (anonymous):

@dpaInc you think you can help me?

OpenStudy (anonymous):

@phi are you able to explain this?

OpenStudy (anonymous):

@ash2326, @UnkleRhaukus anyone able to help?

OpenStudy (saifoo.khan):

Sure. Your question is hard. Everyone can't do it (including me). :/

OpenStudy (anonymous):

ok well i need to go sleep, but going to leave the question open so feel free to answer it anyone. I think the thing i need help with is, from \[\frac{\partial^2 g }{\partial \theta^2} =\frac{d}{d \theta} \left[ \frac{\partial f}{\partial x}\cdot -r\sin\theta +\frac{\partial f}{\partial y}\cdot r\cos\theta\right]\] what \[\frac{d}{d\theta}\left[\frac{\partial f}{\partial x}\right]\] is, so i can carry out the product rule. Please, if you dont mind, tag other helpers in as i will not be able to bump the question

OpenStudy (anonymous):

@hero, @zepdrix , @robtobey ,@Shane_B

OpenStudy (anonymous):

@Wio

OpenStudy (phi):

Have you looked at http://tutorial.math.lamar.edu/Classes/CalcIII/ChainRule.aspx Example 5 near the very bottom (but I would read the whole thing) It looks quite messy.

OpenStudy (anonymous):

@phi thank you for pointing me to that link, its funny as i use that site to help me with other things but missed that part.